OpenStudy (anonymous):

How would you solve the inequality: 36-11x>or=-63

OpenStudy (anonymous):

subtracting 36 from both sides \[-11x \ge -63 -36 \]

OpenStudy (anonymous):

-11x>=-63-36 x<=9

OpenStudy (anonymous):

dividing both sides by -11 \[x \ge -99/-11\]

OpenStudy (anonymous):

\[x \ge 9\]

OpenStudy (anonymous):

sorry made a mistake.... since both sides are negative, we have to change the inequality sign as inik has done

OpenStudy (anonymous):

so final answer is \[x \le9\]

OpenStudy (anonymous):

it's not negative 9?

OpenStudy (anonymous):

no its not -9